OpenStudy (anonymous):

Prove: (secx +cscx) cotx = 1

OpenStudy (aravindg):

Simplify sec x+csc x first!

OpenStudy (anonymous):

I got (cosx/sinx * 1/cosx) + (cosx/sinx * 1/sinx)

OpenStudy (anonymous):

that goes to 1/sinx + cosx/sin^2x

OpenStudy (aravindg):

sec x=.....

OpenStudy (anonymous):

1/cosx

OpenStudy (aravindg):

Then why did you write cos x/sin x?

OpenStudy (anonymous):

cotx= cosx/sinx

OpenStudy (aravindg):

oh wait ..my bad

OpenStudy (aravindg):

dont look at cot x. I told you just simplify sec x+csc x

OpenStudy (anonymous):

1/cosx+1/sinx

OpenStudy (aravindg):

You can do better than that son XD

OpenStudy (anonymous):

cosx+sinx/cosxsinx

OpenStudy (aravindg):

I think your qn is wrong. Just plug in numbers and see. Take x=45 degree. It doesnt hold.

OpenStudy (anonymous):

I don't know, it was given to me like that and I've tried this one before but nothing works...

OpenStudy (aravindg):

Qn is wrong..obviously nothing works.

OpenStudy (anonymous):

I still have more complicated questions to go...sadly

OpenStudy (aravindg):

Np..I can help :)

OpenStudy (anonymous):

1+cosx/1-cosx = (1-cosx)^2/sin^2x so I tried: 1+2cosx+cos^2x/1-cos^2x but I have no clue what I should do next..

mathslover (mathslover):

Good work @AravindG - I am proud of you son! :D And @Greytears - Welcome to OpenStudy

OpenStudy (anonymous):

but sin^2x = 1-cos^2x

OpenStudy (aravindg):

Dude sin^2x should come on NUMERATOR..

OpenStudy (aravindg):

tHANKS @mathslover :)

OpenStudy (anonymous):

I changed it to 1-cos^2x

OpenStudy (aravindg):

1+cos x/1-cos x=sin^2 x/(1-cosx)^2

OpenStudy (aravindg):

sin^2 x=(1-cos x)(1+cos x)

OpenStudy (anonymous):

The problem's: 1+cos x/1-cos x=(1-cosx)^2/sin^2 x

OpenStudy (aravindg):

write sin^2 x=(1-cos x)(1+cos x) and simplify then.

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

oh....